10 babies are born in a hospital on the same day all weigh 2.8 KG each.what would be the standard deviation


1
Expert's answer
2020-09-17T17:07:44-0400
"Var={ \sum x^2 \over n} - \bar {x} ^2""SD = \sqrt {var}"

n=10, x= 2.8

"\sum x^2 = 10 * 2.8^2 = 78.4""\bar {x} = {\sum x \over n} = {10 * 2.8 \over 10} = 7.84"

"{\sum x^2 \over n} = {78.4 \over 10} = 7.84"

"Var = 7.84 - 7.84 =0""SD =\sqrt {var} = \sqrt {0} =0"

Answer: 0